ganeshie8 (ganeshie8):

Find all primes \(p\) such that \(41p+1\) is a perfect square

OpenStudy (fibonaccichick666):

so... 41 is prime ya?

ganeshie8 (ganeshie8):

i bet it is too easy for you to see that lol

OpenStudy (fibonaccichick666):

haha yea, but so isn't this like an extension of mersenne numbers?

ganeshie8 (ganeshie8):

i think this problem is easier than it appears :)

OpenStudy (fibonaccichick666):

oh wait, this is a special number theory prperty isn't it? like \(2^p+1\equiv 0 mod p\)

ganeshie8 (ganeshie8):

we can do it using simple algebra i guess... double click below for a spoiler :P \( \color{white}{41p+1 = n^2 \implies 41p = n^2-1 = (n-1)(n+1)} \)

OpenStudy (fibonaccichick666):

double click?

OpenStudy (fibonaccichick666):

oh you \(\color\white{SUCK}\).

ganeshie8 (ganeshie8):

maybe select the bottom line instead of double clicking..

OpenStudy (fibonaccichick666):

wouldn't that be an equation with 2 unknowns

OpenStudy (fibonaccichick666):

and I got it, try doing it to my "oh you ." post above

ganeshie8 (ganeshie8):

yeah i know i suck

OpenStudy (fibonaccichick666):

:P

ganeshie8 (ganeshie8):

your first reply might help : `so... 41 is prime ya?`

OpenStudy (fibonaccichick666):

I have noo clue... but I need to finish this graph theory homework due in like 8 hours, so I'll look at this some other time

OpenStudy (kainui):

One might be tempted to say 39 and 43... =)

OpenStudy (dan815):

39 aint no prime

OpenStudy (kainui):

It's a hint haha.

OpenStudy (dan815):

oo

OpenStudy (dan815):